OpenStudy (anonymous):

HELP ASAP Scott wants to drive from his house to the video store, which is 5 km away. If he travels at a speed of 50 km/h, how long will the round-trip take him?

OpenStudy (anonymous):

what are you picks?

OpenStudy (anonymous):

@Vodka97

OpenStudy (anonymous):

I Do not have choices.

OpenStudy (anonymous):

Since the distance is 5 km away, and he is going at 50km/h. 5km/(50km/h)= 0.1 hours *60 mins= 6 minutes to get to the store.

OpenStudy (anonymous):

Time = Distance / Speed Time = 5000 /50 * 18/5 time = 2.777 minutes But he's taking a round off trip then the total distance covered is 10 km so Actual time will be. Time = Distance / Speed Time = 10000 /50 * 18/5 time = 5.55 minutes